# Conditioning of Banach Space Valued Gaussian Random Variables: An Approximation Approach Based on Martingales

### v1

@@ D Appendix: Regular Conditional Probabilities
We begin with a general statement that ensures the existence and uniqueness of regular conditional probabilities. For its formulation recall that that a topological space $(T,{\cal T})$ is called Polish, if there exists a metric $\operatorname{d}:T\times T\to[0,\infty)$ generating the topology ${\cal T}$, such that $(T,\operatorname{d})$ is complete and separable. Obviously, separable Banach spaces are Polish spaces. The following theorem, which can be found in e.g. [3, Theorem 44.3], [14, Theorem 10.2.2], and [27, Theorem 8.37], shows that the existence and uniqueness of regular conditional probabilities can be guaranteed as long as the space $T$ in Definition 2.5 is a Polish space and ${\cal B}$ is the corresponding Borel $\sigma$-algebra.
@@
where $\pi_{T_{i}}$ and $\pi_{U_{i}}$ are the coordinate projections for $i=0,1$. Since $\Phi\circ\pi_{T_{0}}$ and $\Psi\circ\pi_{U_{0}}$ are measurable, we then see that $\Upsilon$ is measurable with respect to the product $\sigma$-algebras, see e.g. [27, Corollary 1.82], and in particular our assumption $\mathrm{P}^{(1)}=\mathrm{P}^{(0)}_{(\Phi,\Psi)}=\mathrm{P}^{(0)}_{\Upsilon}$ makes sense. For $B_{1}\in\mathfrak{B}_{1}$ and $A_{1}\in{\cal A}_{1}$ this in turn implies
@@
ii). For $A_{0}\in{\cal A}_{0}$ and $A_{1}\in{\cal A}_{1}$ with $\Psi^{-1}(A_{1})=A_{0}$ and measurable $g:U_{1}\to[0,\infty)$ an application of the transformation formula for integrals of image measures, see e.g. [10, Proposition 2.6.8] yields
@@
$=\int_{U_{1}}\!\boldsymbol{1}_{A_{1}}\cdot g\,\mathrm{d}\mathrm{P}^{(1)}_{U_{1}}\,.$ (D.3)
Applying (D.3) to $g:=\mathrm{P}^{(1)}(B_{1}|\,\cdot\,)$ shows
@@
Our next goal is to investigate Bochner integrals with respect to regular conditional probabilities and how they relate to Banach space valued conditional expectations. Let us begin by introducing the latter. To this end, let $E$ be a separable Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, and ${\cal C}\subset{\cal A}$ be a sub-$\sigma$-algebra. We denote the restriction of $\mathrm{P}$ to ${\cal C}$ by $\mathrm{P}_{|{\cal C}}$ and write accordingly
@@
Let $E$ be a Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, ${\cal C}\subset{\cal A}$ be a sub-$\sigma$-algebra, and $X\in{{\cal L}_{1}(\mathrm{P},E)}$. Then $Y\in{{\cal L}_{1}(\mathrm{P}_{|{\cal C}},E)}$ is called a conditional expectation of $X$ given ${\cal C}$, if
$\int_{\Omega}\boldsymbol{1}_{C}Y\,\mathrm{d}\mathrm{P}_{|{\cal C}}=\int_{\Omega}\boldsymbol{1}_{C}X\,\mathrm{d}\mathrm{P}\,,\qquad\qquad C\in{\cal C}.$
One can show that for each $X\in{{\cal L}_{1}(\mathrm{P},E)}$ there exists a conditional expectation of $X$ given ${\cal C}$, see e.g. [46, Proposition 4.1 in Chapter II.4.1] or [21, Theorem 2.6.20]. Moreover, if we have two conditional expectations $Y_{1}$ and $Y_{2}$ of $X$ given ${\cal C}$, then $\mathrm{P}_{|{\cal C}}(\{Y_{1}\neq Y_{2}\})=0$, see e.g. [13, Corollary 5 in Chapter II.2] or [21, Theorem 2.6.20]. As in the case $E=\mathbb{R}$, we thus write $\mathbb{E}(X|{\cal C})$ for any version of the conditional expectation of $X$ given ${\cal C}$. Here we note that for $X_{1},X_{2}\in{{\cal L}_{1}(\mathrm{P},E)}$ with $\mathrm{P}(\{X_{1}\neq X_{2}\})$ we obviously have $\mathbb{E}(X_{1}|{\cal C})(\omega)=\mathbb{E}(X_{2}|{\cal C})(\omega)$ for $\mathrm{P}_{|{\cal C}}$-almost all $\omega\in\Omega$. Consequently we can view the conditional expectation as a map ${L_{1}(\mathrm{P},E)}\to{L_{1}(\mathrm{P}_{|{\cal C}},E)}$. As usual, in the case ${\cal C}=\sigma(Y)$, we write $\mathbb{E}(X|Y):=\mathbb{E}(X|{\cal C})$.
The next results collects some useful properties of conditional expectations, where we note that most of them are literal generalizations of the well-known case $E=\mathbb{R}$. For a proof we refer to [46, Chapter II.4.1] or [21, Chapter 2.6.d].
@@
Recall that $\mathbb{R}$-valued martingales enjoy well-known convergence results. The following theorem, whose proof can be found in e.g. [46, Theorems 4.1 and 4.2 in Chapter II.4.2] or [21, Theorem 3.3.2 in combination with p. 83], collects analogous results for $E$-valued martingales.
@@
Let $E$ be a Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, and $({\cal C}_{n})$ be a filtration and ${\cal C}_{\infty}:=\sigma({\cal C}_{n}:n\geq 1)$. Moreover, let $1\leq p<\infty$ and $X\in{{\cal L}_{p}(\mathrm{P},E)}$, and for $n\geq 1$ we define
@@
Our next result establishes a disintegration formula for Bochner integrals and regular conditional probabilities. In addition, it shows that regular conditional probabilities can be used to construct conditional expectations. In the case $E=\mathbb{R}$, these results are well-known, see e.g. [14, Theorem 10.2.5] and to some extent also [27, Theorem 8.38], where in all cases actually a more general setting that goes beyond the “product case” is treated.
@@
Proof of Theorem D.6: We first note that for $X:T\times U\to[0,\infty)$ the assertions (D.7) and (D.9) are well-known: Indeed, in this case, the measurability of $Z:U\to[0,\infty]$ defined by (D.8) with $N:=\emptyset$ can be found in e.g. [27, Lemma 14.20]. Moreover, (D.9) is shown in e.g. [27, Theorem 14.29]. For $\mathrm{P}$-integrable $X:T\times U\to[0,\infty)$, the just verified (D.9) then shows that there exists an $N\in{\cal A}$ with $\mathrm{P}_{U}(N)=0$ and $Z(u)\in[0,\infty)$ for all $u\in U\setminus N$, that is, we have (D.7). Finally, if we fix such an $N$ and apply the previous results to $X_{0}:=\boldsymbol{1}_{U\setminus N}X$, we again have the measurability of $Z$ and (D.9).
@@
for some $x_{i}\in E$ and $D_{i}\in{\cal B}\otimes{\cal A}$. For $u\in U$ we have $D_{i,u}:=\{t\in T:(t,u)\in D_{i}\}\in{\cal B}$, see e.g. [27, Lemma 14.13], and hence we see that
$X(\,\cdot\,,u)=\sum_{i=1}^{m}\boldsymbol{1}_{D_{i}}(\,\cdot\,,u)x_{i}=\sum_{i=1}^{m}\boldsymbol{1}_{D_{i,u}}x_{i}$
@@
Since $X_{n}$ is a measurable step function we already know that this definition is possible and that $Z_{n}$ is strongly measurable. Moreover, for each fixed $u$, the sequence $(X_{n}(\,\cdot\,,u))$ consists of measurable step functions approximating $X_{0}(\,\cdot\,,u)$ in the sense of (A.1) and (A.2). By the definition of the Bochner integral (A.4) we thus find
@@
$\int_{T\times U}X_{n}\,\mathrm{d}\mathrm{P}\to\int_{T\times U}X_{0}\,\mathrm{d}\mathrm{P}$ (D.14)
Since we already know that (D.9) holds for all pairs of $X_{n}$ and $Z_{n}$, that is
@@
the just established (D.13) and (D.14) then yield (D.9) in the general case.
@@
for all $u\in U$. From this we can conclude $Z_{A}\circ\pi_{U}=1_{C}\cdot Z\circ\pi_{U}$, and applying (D.9) to $X_{A}$ thus yields
$\int_{T\times U}\boldsymbol{1}_{C}X\,\mathrm{d}\mathrm{P}=\int_{T\times U}X_{A}\,\mathrm{d}\mathrm{P}=\int_{U}Z_{A}\,\mathrm{d}\mathrm{P}_{U}=\int_{U}Z_{A}\,\mathrm{d}\mathrm{P}_{\pi_{U}}=\int_{T\times U}Z_{A}\circ\pi_{U}\,\mathrm{d}\mathrm{P}=\int_{T\times U}\boldsymbol{1}_{C}Z\circ\pi_{U}\,\mathrm{d}\mathrm{P}\,.$

### v3

@@ D Appendix: Regular Conditional Probabilities
We begin with a general statement that ensures the existence and uniqueness of regular conditional probabilities. For its formulation recall that that a topological space $(T,{\cal T})$ is called Polish, if there exists a metric $\operatorname{d}:T\times T\to[0,\infty)$ generating the topology ${\cal T}$, such that $(T,\operatorname{d})$ is complete and separable. Obviously, separable Banach spaces are Polish spaces. The following theorem, which can be found in e.g. [5, Theorem 44.3], [16, Theorem 10.2.2], and [29, Theorem 8.37], shows that the existence and uniqueness of regular conditional probabilities can be guaranteed as long as the space $T$ in Definition 2.5 is a Polish space and ${\cal B}$ is the corresponding Borel $\sigma$-algebra.
@@
where $\pi_{T_{i}}$ and $\pi_{U_{i}}$ are the coordinate projections for $i=0,1$. Since $\Phi\circ\pi_{T_{0}}$ and $\Psi\circ\pi_{U_{0}}$ are measurable, we then see that $\Upsilon$ is measurable with respect to the product $\sigma$-algebras, see e.g. [29, Corollary 1.82], and in particular our assumption $\mathrm{P}^{(1)}=\mathrm{P}^{(0)}_{(\Phi,\Psi)}=\mathrm{P}^{(0)}_{\Upsilon}$ makes sense. For $B_{1}\in\mathfrak{B}_{1}$ and $A_{1}\in{\cal A}_{1}$ this in turn implies
@@
ii). For $A_{0}\in{\cal A}_{0}$ and $A_{1}\in{\cal A}_{1}$ with $\Psi^{-1}(A_{1})=A_{0}$ and measurable $g:U_{1}\to[0,\infty)$ an application of the transformation formula for integrals of image measures, see e.g. [12, Proposition 2.6.8] yields
@@
$=\int_{U_{1}}\!\boldsymbol{1}_{A_{1}}\cdot g\,\mathrm{d}\mathrm{P}^{(1)}_{U_{1}}\,,$ (D.3)
where in the last step we used the already established i). Applying (D.3) to $g:=\mathrm{P}^{(1)}(B_{1}|\,\cdot\,)$ shows
@@
Our next goal is to investigate Bochner integrals with respect to regular conditional probabilities and how they relate to Banach space valued conditional expectations. Let us begin by introducing the latter. To this end, let $E$ be a Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, and ${\cal C}\subset{\cal A}$ be a sub-$\sigma$-algebra. We denote the restriction of $\mathrm{P}$ to ${\cal C}$ by $\mathrm{P}_{|{\cal C}}$ and write accordingly
@@
Let $E$ be a Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, ${\cal C}\subset{\cal A}$ be a sub-$\sigma$-algebra, and $X\in{{\cal L}_{1}(\mathrm{P},E)}$. Then $Z\in{{\cal L}_{1}(\mathrm{P}_{|{\cal C}},E)}$ is called a conditional expectation of $X$ given ${\cal C}$, if
$\int_{\Omega}\boldsymbol{1}_{C}Z\,\mathrm{d}\mathrm{P}_{|{\cal C}}=\int_{\Omega}\boldsymbol{1}_{C}X\,\mathrm{d}\mathrm{P}\,,\qquad\qquad C\in{\cal C}.$
One can show that for each $X\in{{\cal L}_{1}(\mathrm{P},E)}$ there exists a conditional expectation of $X$ given ${\cal C}$, see e.g. [50, Proposition 4.1 in Chapter II.4.1] or [24, Theorem 2.6.20]. Moreover, if we have two conditional expectations $Z_{1}$ and $Z_{2}$ of $X$ given ${\cal C}$, then $\mathrm{P}_{|{\cal C}}(\{Z_{1}\neq Z_{2}\})=0$, see e.g. [15, Corollary 5 in Chapter II.2] or [24, Theorem 2.6.20]. As in the case $E=\mathbb{R}$, we thus write $\mathbb{E}(X|{\cal C})$ for any version of the conditional expectation of $X$ given ${\cal C}$. Here we note that for $X_{1},X_{2}\in{{\cal L}_{1}(\mathrm{P},E)}$ with $\mathrm{P}(\{X_{1}\neq X_{2}\})=0$ we obviously have $\mathbb{E}(X_{1}|{\cal C})(\omega)=\mathbb{E}(X_{2}|{\cal C})(\omega)$ for $\mathrm{P}_{|{\cal C}}$-almost all $\omega\in\Omega$. Consequently we can view the conditional expectation as a map ${L_{1}(\mathrm{P},E)}\to{L_{1}(\mathrm{P}_{|{\cal C}},E)}$. As usual, in the case ${\cal C}=\sigma(Y)$, we write $\mathbb{E}(X|Y):=\mathbb{E}(X|{\cal C})$.
The next results collects some useful properties of conditional expectations, where we note that most of them are literal generalizations of well-known properties in the case $E=\mathbb{R}$. For a proof we refer to [50, Chapter II.4.1] or [24, Chapter 2.6.d].
@@
Recall that $\mathbb{R}$-valued martingales enjoy well-known convergence results. The following theorem, whose proof can be found in e.g. [50, Theorems 4.1 and 4.2 in Chapter II.4.2] or [24, Theorem 3.3.2], collects analogous results for $E$-valued martingales.
@@
Let $E$ be a Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $({\cal C}_{n})$ be a filtration, and ${\cal C}_{\infty}:=\sigma({\cal C}_{n}:n\geq 1)$. Moreover, let $1\leq p<\infty$ and $X\in{{\cal L}_{p}(\mathrm{P},E)}$, and for $n\geq 1$ we define
@@
Our next result establishes a disintegration formula for Bochner integrals and regular conditional probabilities. In addition, it shows that regular conditional probabilities can be used to construct conditional expectations. In the case $E=\mathbb{R}$, these results are well-known, see e.g. [16, Theorem 10.2.5] and to some extent also [29, Theorem 8.38], where in all cases actually a more general setting that goes beyond the “product case” is treated.
@@
Proof of Theorem D.6: We first note that for $X:T\times U\to[0,\infty)$ the assertions (D.7) and (D.9) are well-known: Indeed, in this case, the measurability of $Z:U\to[0,\infty]$ defined by (D.8) with $N:=\emptyset$ can be found in e.g. [29, Lemma 14.20]. Moreover, (D.9) is shown in e.g. [29, Theorem 14.29]. For $\mathrm{P}$-integrable $X:T\times U\to[0,\infty)$, the just verified (D.9) then shows that there exists an $N\in{\cal A}$ with $\mathrm{P}_{U}(N)=0$ and $Z(u)\in[0,\infty)$ for all $u\in U\setminus N$, that is, we have (D.7). Finally, if we fix such an $N$ and apply the previous results to $X_{0}:=\boldsymbol{1}_{U\setminus N}X$, we again have the measurability of $Z$ and (D.9).
@@
for some $x_{i}\in E$ and $D_{i}\in{\cal B}\otimes{\cal A}$. For $u\in U$ we have $D_{i,u}:=\{t\in T:(t,u)\in D_{i}\}\in{\cal B}$, see e.g. [29, Lemma 14.13], and hence we see that
$X(\,\cdot\,,u)=\sum_{i=1}^{m}\boldsymbol{1}_{D_{i}}(\,\cdot\,,u)x_{i}=\sum_{i=1}^{m}\boldsymbol{1}_{D_{i,u}}x_{i}$ (D.10)
@@
Since $X_{n}$ is a measurable step function we already know that this definition is possible and that $Z_{n}$ is strongly measurable. Moreover, for each fixed $u$, the sequence $(X_{n}(\,\cdot\,,u))$ consists of measurable step functions, see (D.10), which approximates $X_{0}(\,\cdot\,,u)$ in the sense of (A.1) and (A.2). By the definition of the Bochner integral (A.4) we thus find
@@
$\int_{T\times U}X_{n}\,\mathrm{d}\mathrm{P}\to\int_{T\times U}X_{0}\,\mathrm{d}\mathrm{P}\,.$ (D.15)
Moreover, we already know that (D.9) holds for all pairs of $X_{n}$ and $Z_{n}$, that is
@@
and therefore the just established (D.14) and (D.15) then yield (D.9) in the general case.
@@
for all $u\in U$. From this we can conclude $Z_{A}\circ\pi_{U}=1_{C}\cdot(Z\circ\pi_{U})$, and applying (D.9) to $X_{A}$ thus yields
$\int_{T\times U}\boldsymbol{1}_{C}X\,\mathrm{d}\mathrm{P}=\int_{T\times U}X_{A}\,\mathrm{d}\mathrm{P}=\int_{U}Z_{A}\,\mathrm{d}\mathrm{P}_{U}=\int_{U}Z_{A}\,\mathrm{d}\mathrm{P}_{\pi_{U}}$ $=\int_{T\times U}Z_{A}\circ\pi_{U}\,\mathrm{d}\mathrm{P}$
$=\int_{T\times U}\boldsymbol{1}_{C}\cdot(Z\circ\pi_{U})\,\mathrm{d}\mathrm{P}\,.$